OpenStudy (anonymous):

Find h'(1) given that f(1) = -6, f '(1) = -5, g(1) = 5, and g'(1) = 6. (a) h(x) = 3f(x) - 3g(x)

OpenStudy (anonymous):

h'(x) = 3 f'(x) - 3 g'(x) h'(1) = 3 (-5) -3 (6) = -15-18 = -33